How much energy in kWh would a heater rated at 7 kW use when run for 9000 seconds? If the cost of electricity is 1.4 $/kWh, how much does it cost to run the heating system for 9000 seconds ?
Answer #1

Time 9000 seconds

9000/3600= 2.5 hours

Electricity consumed =7*2.5= 17.5 KWh

Cost =17.5*1.4= 24.5 dollars Ans